# Theoretical Bounds for Stable In-Context Learning

### v2

@@ Abstract
Figure 1: Schematic comparison of ICL prompt-length selection. (a) Conventional trial-and-error or heuristic approaches leave uncertainty about the minimal effective number of demonstrations (K). (b) Our proposed spectral-coverage method provides a theoretically grounded, observable proxy to determine a stable minimum prompt length (K*), ensuring reliable in-context learning.
@@ 1 Introduction
In-context learning (ICL) has become a key paradigm for large language models to adapt to downstream tasks, as it enables task inference without parameter updates, solely through prompting(Wei et al., 2022; Tanwar et al., 2023; Dong et al., 2024). However, the output of ICL is sensitive to the selection and order of specific examples in the prompt(Xie et al., 2022; Zhao et al., 2021; Chen et al., 2023; Li and Qiu, 2023; Fang et al., 2025), making its stability difficult to guarantee. An unresolved core question is: how many examples are needed to ensure stable ICL predictions? The lack of a computable answer to this question leads to reliance on heuristics or excessively long prompts in practice, sacrificing efficiency and failing to provide reliability guarantees.
This paper studies the stability of ICL and the minimal necessary prompt length. We formally define ICL stability in a probabilistic sense: for the same query, two independently sampled sets of examples should induce similar model output distributions. Directly evaluating this stability requires extensive repeated queries to the model, which is costly. To address this, we introduce an observable proxy objective: spectral coverage of example representations, specifically manifested as the minimum eigenvalue of the regularized empirical second-moment (uncentered covariance) matrix. Our goal is to estimate, from data, the minimum number of examples $K^{*}$ required to meet this proxy objective under given stability tolerances $(\tau,\xi)$, and to provide non-asymptotic theoretical guarantees.
@@
Theory (sufficient condition & threshold interpretation): We derive a non-asymptotic sufficient condition for spectral coverage under sub-Gaussian representations, expressed as an explicit lower bound on the required prompt length $K$. Interpreting this condition yields a conservative upper bound on the (unknown) stability threshold, and we further analyze robustness under representation drift, heavy tails, and weak dependence. Observable estimator (no population priors): We design a fully observable two-stage procedure that estimates the required spectral statistics from a small pilot prompt and returns a concrete prompt length with a prescribed failure probability.
Empirical alignment (knee, stability diagnostics, calibration): Across datasets, task families, and commercial API models, we verify that the predicted thresholds upper-bound empirical knee-points, and that a lightweight validation-only calibration (global scale and quantile level) substantially tightens the gap while preserving conservative ordering.
@@ 2 Related Works
Theoretical foundations of ICL. A growing body of work formalizes ICL within established learning frameworks. PAC-style analyses establish learnability conditions under mixture models, while other studies treat Transformers as algorithms with task-dependent sample complexity(Frei and Vardi, 2024; Magen and Vardi, 2025; Li et al., 2025). Asymptotic analyses, particularly in linearized settings, describe phase transitions in learning curves, and task-specific results for regression and structured prediction further refine these analyses(Chang et al., 2025; Takanami et al., 2025). While these provide key asymptotic insights, they often lack non-asymptotic, quantitative guarantees that prescribe a concrete number of demonstrations $K$ for a given task and model—a critical need for practical prompt design. Our work fills this gap by deriving a non-asymptotic lower bound on $K$ based on spectral coverage of demonstration representations, offering a certificate of stability beyond mere eventual convergence.
@@
Mechanistic and empirical perspectives. Understanding how ICL emerges internally has been explored through mechanistic studies linking Transformer internal representations to ICL behavior(Li and Qiu, 2023; Fang et al., 2025). Empirically, comprehensive benchmarks and toolkits have standardized evaluation pipelines(Chen et al., 2025; Rieff et al., 2025). Notably, independent empirical work observes that performance knee-points often require far fewer demonstrations than heuristic rules suggest(Liu et al., 2022; Hendel et al., 2023). Our theory provides a direct mechanistic and quantitative explanation for this phenomenon: knee-points coincide with the moment when the empirical second-moment (uncentered covariance) matrix of demonstrations achieves sufficient spectral coverage, a condition we formalize and make computable. While these empirical and mechanistic studies document phenomena and mechanisms, our analysis provides a predictive why and when from a statistical learning perspective.
Our positioning. In summary, prior theoretical work offers asymptotic characterizations rather than practical lower bounds; calibration research optimizes outputs for fixed prompts rather than determining prompt size; empirical/mechanistic studies describe phenomena that we can now quantitatively explain. Our contribution complements these lines by introducing a spectral coverage proxy for ICL stability, deriving its non-asymptotic lower bound, and providing an efficient two-stage algorithm to estimate it from data. This approach offers a missing normative, sample-complexity-oriented perspective, turning qualitative observations about knee-points into a quantitatively verifiable and actionable stability condition.
@@ 3.1 Problem setup and proxy objective
Definition A.1 formalizes stability in terms of the agreement between output distributions induced by two independently sampled prompts of the same length. Concretely, consider a classification task where each prompt consists of $K$ labeled demonstrations. Two prompts $P_{K}$ and $P^{\prime}_{K}$ are formed by independently sampling and ordering demonstrations from the same task pool. If the model is unstable at length $K$, small changes in the selected demonstrations may lead to noticeably different output distributions—for example, flipping the predicted label or substantially redistributing probability mass among competing classes. In contrast, stability implies that, despite variation in the demonstrations, the model produces highly similar predictive distributions, reflecting insensitivity to the specific realization of the prompt. This notion captures robustness to prompt composition rather than average accuracy, and is task-agnostic.
@@
Definition A.1 is stated for a general LLM output distribution. Our theory does not model the full generator; instead, it analyzes a computable spectral-coverage proxy under a frozen representation map $\phi$ and standard concentration assumptions. In black-box API settings, the generator typically does not expose hidden states, so in practice $\phi$ can be instantiated using a fixed external encoder. In our experiments, $\phi(x)$ is instantiated as a fixed sentence-level embedding $e(x)$ of the input text, and we form $\widehat{\Sigma}_{K}:=\frac{1}{K}\sum_{i=1}^{K}e(x_{i})e(x_{i})^{\top}$ from the selected demonstrations, denoting the empirical second-moment (uncentered covariance) matrix of the representation vectors induced by the $K$ demonstrations. (see Appendix C for encoder and preprocessing details). All finite-sample guarantees in this section concern the resulting proxy quantities computed in the $\phi$-space.
Directly verifying Definition A.1 is expensive because it requires repeated model queries under many resampled prompts. We therefore study an observable sufficient condition based on the spectral coverage of demonstration representations. Let $\phi:\mathcal{X}\to\mathbb{R}^{d}$ be a fixed feature map and define the (uncentered) population second-moment matrix
@@ 3.5 Calibration and practical recommendations
The bound in Proposition 3.3 and the plug-in rule (8) are intentionally conservative: they protect against worst-case directions in the feature space and use high-probability concentration. In practice, two issues commonly arise: (i) the smallest eigenvalue is overly sensitive to noise in near-null directions, and (ii) constants hidden in non-asymptotic inequalities can inflate the suggested prompt length. This section summarizes lightweight calibration strategies that preserve the monotonic behavior of the proxy while producing practically useful $K$.
@@
Default choices.
@@ 4.1 Experimental setup
4.2 Measuring the Knee-Point
Although $\widehat{K}_{\mathrm{knee}}$ is defined via accuracy saturation rather than Definition A.1, we use it only as an empirical baseline for comparison. To support the stability interpretation, we also directly measure distributional stability under repeated demonstration resampling on a held-out subset; see Appendix L.
We summarize prediction quality using the error ratios
where $K^{*}$ denotes the proxy-based prompt-length estimate produced by our observable estimator (Algorithm 1), and $\widehat{K}_{\mathrm{cal}}$ applies validation-only calibration as in (12) (Section 3.5).
4.3 Numerical sensitivity of spectral proxies
Protocol.
@@ 5 Conclusion
This work provides a theoretically grounded, computable sufficient sample-size requirement for stable In-Context Learning (ICL) under a spectral-coverage proxy, together with an observable two-stage estimator that returns a concrete prompt length with a prescribed failure probability. By linking ICL stability trends to measurable high-dimensional spectral statistics, our analysis bridges statistical theory and practical prompt design, reducing unnecessary long-context overhead while making reliability claims verifiable. Future work may relax the assumption of fixed representations, refine calibration under more dynamic settings, and extend the framework to multi-modal ICL.
@@ Appendix A Complete Theory Chain
Definition A.1 is stated for a general LLM output distribution. Our theoretical analysis does not model the full LLM; instead it analyzes a computable spectral-coverage proxy under a frozen representation map $\phi$ and standard sub-Gaussian concentration.
@@
In particular, when we use the proxy to predict the generator’s ICL stability (Definition A.1), we rely on an empirical proxy–generator alignment premise, rather than claiming a first-principles equivalence. We validate this premise by direct distributional stability measurements and comparisons in Appendix L.
@@
Directly verifying Definition A.1 is expensive because it requires repeated model queries under many resampled prompts. We therefore study an observable sufficient condition based on the spectral coverage of demonstration representations. Concretely, we form an empirical second-moment matrix $\widehat{\Sigma}_{K}$ from representations of the $K$ demonstrations and require its spectrum to be bounded away from degeneracy. Sections 3–4 provide non-asymptotic bounds and an observable two-stage estimator for this proxy, while later experiments validate that the proxy upper-bounds empirical knee-points and can be calibrated.
@@ I.10 Model Scale
The proxy objective in (2) is parameterized by a target spectral floor $\delta$, which controls how well-conditioned the demonstration second moment must be to certify the surrogate stability event. In Algorithm 1, the check $\widehat{\Delta}=\underline{\lambda}-\delta>0$ is necessary; if the algorithm returns infeasible target, it means that the requested $\delta$ is larger than the estimated population floor at the desired confidence level, under the current feature map and ridge level. In practice we recommend selecting $\delta$ by validation: sweep $\delta$ on a small held-out set, choose the smallest value that achieves the desired stability/accuracy trade-off, and then lock it for test-time estimation. When available, Appendix L provides a direct estimator for distributional stability in the sense of Definition A.1 that can serve as the validation signal. If infeasibility persists, reduce $\delta$, increase the ridge via $\epsilon_{\rho}$, or use the quantile-floor variant described in Section 3.5.

### v3

# Stability of In-Context Learning: A Spectral Coverage Perspective

@@ Abstract
Figure 1: Schematic comparison of ICL prompt-length selection. (a) Conventional trial-and-error or heuristic approaches leave uncertainty about the minimal effective number of demonstrations (K). (b) Our proposed spectral-coverage method provides a theoretically grounded, observable proxy to select a conservative prompt length ($K^{\star}$) that targets distributional stability under demonstration resampling (Definition 3.1).
@@ 1 Introduction
In-context learning (ICL) has become a key paradigm for large language models to adapt to downstream tasks, as it enables task inference without parameter updates, solely through prompting(Wei et al., 2022; Tanwar et al., 2023; Dong et al., 2024). However, the output of ICL is sensitive to the selection and order of specific examples in the prompt(Xie et al., 2022; Zhao et al., 2021; Chen et al., 2023; Li and Qiu, 2023; Fang et al., 2025), making its stability difficult to guarantee. An unresolved core question is: how can we select a minimal prompt length that yields reliable ICL behavior, when distributional stability under demonstration resampling is costly to measure directly? The lack of a computable answer to this question leads to reliance on heuristics or excessively long prompts in practice, sacrificing efficiency and failing to provide reliability guarantees.
This paper studies the stability of ICL and the minimal necessary prompt length. We formally define ICL stability in a probabilistic sense: for the same query, two independently sampled sets of examples should induce similar model output distributions. Directly evaluating this stability requires extensive repeated queries to the model, which is costly. Throughout, we treat distributional stability under demonstration resampling as the conceptual target. However, directly estimating it requires repeated model queries for many independently resampled prompts, which is prohibitive for large-scale benchmarking. Our theoretical results therefore concern an observable spectral-coverage proxy computed from demonstration representations. Empirically, we evaluate this proxy at scale using accuracy knee-points only as a low-cost surrogate for the prompt-length transition, and we separately validate the proxy–stability connection via direct resampling-based distributional measurements on a smaller subset (Appendix L). To address this, we introduce an observable proxy objective: spectral coverage of example representations, specifically manifested as the minimum eigenvalue of the regularized empirical second-moment (uncentered covariance) matrix. Our goal is to estimate, from data(Chang and Jia, 2023), the minimum number of examples $K^{*}$ required to meet this proxy objective under given stability tolerances $(\tau,\xi)$, and to provide non-asymptotic theoretical guarantees.
@@
Theory (computable proxy with non-asymptotic guarantees): We derive a non-asymptotic sufficient condition for achieving a spectral-coverage proxy under sub-Gaussian representations, expressed as an explicit lower bound on the required prompt length $K$. This yields a conservative, fully observable prompt-length recommendation with a prescribed failure probability. We further analyze robustness under representation drift, heavy tails, and weak dependence.
Observable estimator (no population priors): We design a fully observable two-stage procedure that estimates the required spectral statistics from a small pilot prompt and returns a concrete prompt length with a prescribed failure probability for the proxy event.
Empirical validation (surrogates at scale, stability checks on a subset): Across datasets, task families, and commercial API models, we show that the proxy-based prompt-length recommendations conservatively upper-bound empirical accuracy knee-points, which we use only as a scalable surrogate for the prompt-length transition. We additionally perform direct resampling-based distributional stability measurements on a smaller held-out subset to validate the intended stability interpretation (Appendix L). A lightweight validation-only calibration (global scale and quantile level) further tightens the conservatism while preserving conservative ordering.
@@ 2 Related Works
Theoretical foundations of ICL. A growing body of work formalizes ICL within established learning frameworks. PAC-style analyses establish learnability under mixture models, and other studies treat Transformers as algorithms with task-dependent sample complexity(Frei and Vardi, 2024; Magen and Vardi, 2025; Li et al., 2025). Linearized asymptotic analyses characterize phase transitions, with task-specific results for regression and structured prediction(Chang et al., 2025; Takanami et al., 2025). Relatedly, spectrum-oriented post-training interventions aim to improve distributional coverage and in-context steerability(Sorensen et al., 2025). However, these lines generally do not provide non-asymptotic guarantees that yield a concrete demonstration length $K$ for practical prompt design. We address this by deriving a non-asymptotic lower bound on $K$ from spectral coverage of demonstration representations, yielding a computable stability proxy.
@@
Mechanistic and empirical perspectives. Understanding how ICL emerges internally has been explored through mechanistic studies linking Transformer representations to ICL behavior(Li and Qiu, 2023; Fang et al., 2025). Empirically, benchmarks and toolkits standardize evaluation(Chen et al., 2025; Rieff et al., 2025), and independent work finds that performance knee-points often require far fewer demonstrations than heuristics suggest(Liu et al., 2022; Hendel et al., 2023). Relatedly, in-context clustering work analyzes geometric structure in attention/representations and leverages spectral clustering-style tools in LLM-driven ICL settings(Wang et al., 2025). Complementary to resampling-driven instability, latent shifts have also been studied as a source of ICL unreliability, with self-training proposed as a stabilization mechanism(Jukić and Šnajder, 2024). Our theory provides a mechanistic explanation for why accuracy knee-points often appear near the onset of sufficient spectral coverage of demonstration representations, a condition we formalize and make computable.
@@ 3.1 Problem setup and proxy objective
Definition 3.1 formalizes stability in terms of the agreement between output distributions induced by two independently sampled prompts of the same length. Concretely, consider a classification task where each prompt consists of $K$ labeled demonstrations. Two prompts $P_{K}$ and $P^{\prime}_{K}$ are formed by independently sampling and ordering demonstrations from the same task pool. If the model is unstable at length $K$, small changes in the selected demonstrations may lead to noticeably different output distributions—for example, flipping the predicted label or substantially redistributing probability mass among competing classes. In contrast, stability implies that, despite variation in the demonstrations, the model produces highly similar predictive distributions, reflecting insensitivity to the specific realization of the prompt. This notion captures robustness to prompt composition rather than average accuracy, and is task-agnostic.
@@
Definition 3.1 concerns the LLM output distribution, but our analysis uses an observable proxy computed from a representation map $\phi:\mathcal{X}\to\mathbb{R}^{d}$. In black-box API settings we instantiate $\phi$ with a frozen external encoder and compute all spectral quantities in this encoder space; implementation details are in Appendix C.
Directly verifying Definition 3.1 is expensive because it requires repeated model queries under many resampled prompts. We therefore study an observable sufficient condition based on the spectral coverage of demonstration representations. Let $\phi:\mathcal{X}\to\mathbb{R}^{d}$ be a fixed feature map and define the (uncentered) population second-moment matrix
@@ 3.5 Calibration and practical recommendations
The bound in Proposition 3.3 and the plug-in rule (8) are intentionally conservative: they protect against worst-case directions in the feature space and use high-probability concentration. Importantly, calibration is introduced as a practical correction for proxy–generator mismatch and finite-sample conservatism; it does not change the non-asymptotic guarantee for the uncalibrated proxy bound, and should be viewed as an empirical tightening step rather than a first-principles refinement of the theory. In practice, two issues commonly arise: (i) the smallest eigenvalue is overly sensitive to noise in near-null directions, and (ii) constants hidden in non-asymptotic inequalities can inflate the suggested prompt length. This section summarizes lightweight calibration strategies that preserve the monotonic behavior of the proxy while producing practically useful $K$.
@@
Default practical choices.
Robust default settings (ridge level, spectral threshold choice, and failure-budget split) are summarized in Appendix J.5.
@@ 4.1 Experimental setup
4.2 Measuring the Knee-Point
Although $\widehat{K}_{\mathrm{knee}}$ is defined via accuracy saturation rather than Definition 3.1, we use it only as an empirical baseline for comparison. A plateau in accuracy alone does not certify distributional stability, particularly under stringent stability tolerances or open-generation settings. To support the intended stability interpretation, we additionally measure distributional stability under repeated demonstration resampling on a held-out subset; see Appendix L.
We summarize prediction quality using the error ratios
where $K^{*}$ denotes the proxy-based prompt-length estimate produced by our observable estimator (Algorithm 1), and $\widehat{K}_{\mathrm{cal}}$ applies validation-only calibration as in (12) .
4.3 Numerical sensitivity of spectral proxies
Protocol.
@@ 5 Conclusion
This work provides a theoretically grounded, computable sufficient sample-size requirement for ICL under a spectral-coverage proxy, together with an observable two-stage estimator that returns a concrete prompt length with a prescribed failure probability. By linking ICL stability trends to measurable high-dimensional spectral statistics, our analysis bridges statistical theory and practical prompt design, reducing unnecessary long-context overhead while making reliability claims verifiable. Future work may relax the assumption of fixed representations, refine calibration under more dynamic settings, and extend the framework to multi-modal ICL(Oorloff et al., 2025).
@@ Appendix A Complete Theory Chain
Definition 3.1 is stated for a general LLM output distribution. Our theoretical analysis does not model the full LLM; instead it analyzes a computable spectral-coverage proxy under a frozen representation map $\phi$ and standard sub-Gaussian concentration.
@@
In particular, when we use the proxy to predict the generator’s ICL stability (Definition 3.1), we rely on an empirical proxy–generator alignment premise, rather than claiming a first-principles equivalence. We validate this premise by direct distributional stability measurements and comparisons in Appendix L.
@@
Directly verifying Definition 3.1 is expensive because it requires repeated model queries under many resampled prompts. We therefore study an observable sufficient condition based on the spectral coverage of demonstration representations. Concretely, we form an empirical second-moment matrix $\widehat{\Sigma}_{K}$ from representations of the $K$ demonstrations and require its spectrum to be bounded away from degeneracy. Sections 3–4 provide non-asymptotic bounds and an observable two-stage estimator for this proxy, while later experiments validate that the proxy upper-bounds empirical knee-points and can be calibrated.
@@ I.10 Model Scale
The proxy objective in (2) is parameterized by a target spectral floor $\delta$, which controls how well-conditioned the demonstration second moment must be to certify the surrogate stability event. In Algorithm 1, the check $\widehat{\Delta}=\underline{\lambda}-\delta>0$ is necessary; if the algorithm returns infeasible target, it means that the requested $\delta$ is larger than the estimated population floor at the desired confidence level, under the current feature map and ridge level. In practice we recommend selecting $\delta$ by validation: sweep $\delta$ on a small held-out set, choose the smallest value that achieves the desired stability/accuracy trade-off, and then lock it for test-time estimation. When available, Appendix L provides a direct estimator for distributional stability in the sense of Definition 3.1 that can serve as the validation signal. If infeasibility persists, reduce $\delta$, increase the ridge via $\epsilon_{\rho}$, or use the quantile-floor variant described in Section 3.5.